OpenStudy (anonymous):

12.given that the equation of a circle is x^2+y^2-10x+4y+13=0,find its cente and its raius?

OpenStudy (anonymous):

complete the squares as follows (x-5)^2 + (y+2)^2 = -13+25+4 = 16 Center: (5, -2) Radius = 4